In a recent tomato eating contest, the average number of tomatoes a person ate was 21 witha standard deviation of 4. Find the Z- score of a person who ate the following amount of tomatoes

A) 27 Tomatoes

B) 19 Tomatoes

C) 21 Tomatoes

Using the information above, find the corresponding number of tomatoes eaten for the given Z-score:

A) Z= 2.1

B) Z= -0.9

C) Z = 1
Answer #1

Ans:

i)

a)

z=(27-21)/4=1.5

b)

z=(19-21)/4=-0.5

c)

z=(21-21)/4=0

ii)

a)z=21+2.1*4=29.4

b)z=21-0.9*4=17.4

c)z=21+1*4=25